# Context-Aware Code Generation Framework for Code Repositories: Local, Global, and Third-Party Library Awareness

### 03 Mar 2024 v3

@@ I Introduction
In the realm of Software Engineering, the advent of Large Language Models (LLMs) has revolutionized various tasks, including code generation [1, 2, 3, 4, 5]. The automatic generation of code by these models significantly reduces development time and effort [6]. Specifically, in the process of evolving software projects, software engineering developers are frequently tasked to add new features and functionality [7]. LLMs such as ChatGPT [8], GitHub Copilot [9], Codex [10] are capable of automatically generating code based on the developers’ requirements (e.g., natural language descriptions of functions and function definitions).
A significant gap still persists between the code generated by LLMs and that crafted by human developers. When adding new functionality to a specific module, developers already had sufficient knowledge about the current task and the entire code repository. Also, according to a previous study [11], developers often expect that code generation tools could be aware of more context/knowledge by using specific frameworks/libraries [12]. In addition, they hope the generated code maximum utilizes specific third-party libraries (e.g., invocation of an API in a library) [13, 14, 15]. Above all, in contrast to straightforward LLM-based code generation, developers especially consider three primary categories of information:
@@
Global Information: This category involves functions from other code files within the same repository, accessible via fully qualified import API names. A previous research highlighted that only around 30% of methods operate independently in open-source projects [16]. Hence, these external functions are essential for implementing new functions without duplicating code. For example, developers import functions “from unstructured.partition.text_type import is_bulleted_text” from other modules, instead of writing a function to judge “bulleted text” from scratch.
@@
Unfortunately, none of existing code generation methods take above three information into consideration simultaneously. As the most common code generation tool, ChatGPT [8] demonstrates commendable coding ability [17] but relies solely on user input. GitHub Copilot [9], which explicitly utilizes local information spanning the current working context, often struggles to harness the full potential of global information presented in the entire code repository [^1] https://githubnext.com/projects/copilot-view/. When these methods disregard local and global module information, they risk omitting crucial logic during code generation [18]. For instance, they might forget to extract text from a tag before attempting to identify it as a bullet point (is_bulleted_text). Furthermore, neglecting third-party library information may cause the LLMs to randomly select third-party libraries and generate code based on them. If the selected libraries third-party are unsupported under the given environment, it can lead to compatibility issues, such as “ModuleNotFoundError: No module named xxx.
Therefore, to encounter the insufficient information awareness issue in code generation tasks, we introduce novel $\mathbf{A^{3}}$-CodGen (Local Aware, Global Aware, and Third-Party-Library Aware). This framework systematically mines and utilizes external knowledge from the local module, global module, and third-party libraries. First, our framework begins by acquiring foundational information about functions and third-party libraries within the code repository. This initial step provides a solid function base and a third-party-library base that supports subsequent information retrieval processes. Next, it conducts knowledge retrieval from the function base, the third-party library base, and the current file. These retrieved insights are systematically organized into three distinct categories of code repository-aware knowledge. Subsequently, the framework seamlessly fuses these three types of code repository-aware knowledge into a cohesive prompt. This prompt serves as the input to the foundational LLM, enabling the model to generate code that closely resembles the work of human developers.
@@
  • We introduce a novel framework, dubbed $\mathbf{A^{3}}$-CodGen Framework, effectively knowledge mining and knowledge utilization, aiding models in better understanding and fully utilizing the knowledge within code repositories.
@@
  • We open-source our replication package [19], including the dataset, source code, and experimental results, for follow-up research.
@@ II-A Limitation of Existing Code Generation Approaches
Benefiting from the rapid development of Transformer structures and pre-trained LLMs in the early 2020s, code-LLMs experienced significant growth. Transformer code models, exemplified by milestone works such as CodeT5 [20], CodeBERT [21], and GraphCodeBERT [22], dominated the scene. More recently, decoder-only Transformer models have become the mainstream in generative models. Represented by Codex [23], CodeGen [24], and others like [25, 26, 27, 28, 29].
Nowadays, to Software Engineering developers, LLM-empowered code generation has become an irreplaceable assistant tool in their development cycle [30, 31]. Although the code-LLM has been developed for years, those code generation tools still play a supporting role in the actual SE development. Essentially, there are significant differences between the code generated by LLMs and that crafted by human developers. Specifically, we conclude the phenomenon in the following threefold:
@@
    Discussion 1: Providing longer prompts with richer content seems a plausible solution. With the growth of LLM size, the maximum input length increases from 512 tokens for the early models [32] to 32,000 tokens for the extended GPT-4 models [33]. However, it is still impossible to feed the entire code repository and all other relevant information into the LLMs. Furthermore, studies have demonstrated that irrelevant context in prompts not only affects model efficiency and increases expenses, but also potentially compromises the model performance [34]. Therefore, to maintain the practicability and effectiveness simultaneously for code generation tasks, external context information should be extracted and selectively fed into LLMs.
  • Observation 2: Existing methods focus on generating independent and often small-scale code units [35, 16]. According to a previous study [35], the code generated by LLMs shows lower method-level correctness. Although various benchmarks have been proposed (Details are discussed in the Related Work (Section VII)), newly generated code still remains questionable.
@@
  • Observation 3: Some code-LLMs (e.g., GPT-4) have developed superior skills in simple algorithms such as greedy algorithm and searching strategies, but they perform poorly in complicated data structures such as tree structure and graph theory [36].
@@ II-B Code Generation in Code Repository
During software development, programmers typically work within a specific code environment, essentially a designated code repository. This repository encompasses all project modules, functions in these modules, and third-party libraries, serving as the foundational framework for building and expanding software applications. In this context, developers, especially experienced ones, must thoroughly navigate and understand this code repository environment to ensure seamless integration of newly added code with the existing one. When it comes to LLM-based code generation techniques, solely relying on inputting developer requirements to these models fails to fully utilize the potential of LLMs in interacting with existing code repository environments. Specifically, these techniques are blind to the repository’s unique naming conventions and programming styles, particularly when the completion target is relevant to cross-file context [37, 38, 39]. Instead of using a single code file, well-designed programs emphasize separating complicated functionality into independent modules, each implementing only one aspect of the desired functionality [40, 41].
@@
Furthermore, since the model is unaware of the existence of the “check text for bullet characters” function in the global module, it defines a new function is_bullet_text (highlighted in blue) in the generated code. A large part of this code duplicates the function is_bullet_text in the global module unstructured/partition/text_type.py, resulting in code redundancy. It also brings additional concerns about the correctness of the newly generated function.
Additionally, since the model is unaware of which third-party libraries are available, it arbitrarily imports and uses the bs4 (BeautifulSoup4) library. This can result in the generated code that refers to a library that is not available, leading to the bug “ModuleNotFoundError: No module named ***”.
@@ II-C Bridging the Gap: Making LLMs Aware of Three Types of Repository Knowledge
As shown in Fig. 1 Local Module Aware & Global Module Aware & Third-Party Library Aware Code Generation, when the developer requirement is entered, since the model is aware of the knowledge of the local module, the generated code reuses the _construct_text function (highlighted in yellow) in the local module to extract the text content in the text label element. In addition, because the model is aware of knowledge of the global module, the generated code reuses the is_bulleted_text function (highlighted in blue) from the global module unstructured/partition/text_type.py, thus reducing redundant code definitions. In addition, the model is aware of the existence of third-party libraries and therefore reuses the lxml library to ensure its availability.
@@ III $A^{3}$odGen Framework
In this section, we elaborate on our novel $A^{3}$-CodGen Framework to harness the full potential of code repository knowledge for assisting LLMs in code generation. As depicted in Fig. 2, the $A^{3}$-CodGen Framework comprises three essential phases. First, we obtain the function base and third-party library base (Section III-A), providing sufficient knowledge to support subsequent retrieval. Then, we retrieve the code repository knowledge from other code files, the third-party library repository, and the current file. After that, we organize them into 3 types of code repository aware knowledge (Section III-B). Finally, the three code repository-aware knowledge are fused into a prompt (Section III-C), serving as input to the foundation LLM, to ultimately generate codes with three essential awareness.
@@ III-A Repository Knowledge Base Construction
Code repositories contain a wealth of knowledge to exploit, but this knowledge usually exists in an unstructured form and cannot be directly utilized effectively, scattering in the code repository. Therefore, we need to perform systematic knowledge extraction to transform and compress this knowledge into a structured form in order to build a knowledge base to support subsequent knowledge utilization, such as knowledge retrieval. Thus, we perform code repository information extraction (Section III-A1) and third-party library extraction (Section III-A2) to obtain the function base and third-party-library base, respectively.
@@ III-A1 Code Repository Information Extraction
Basic Code Knowledge Extraction. Abstract Syntax Tree (AST) is a tree representation of the source code, and studies have demonstrated its usefulness on various SE tasks [42]. In this study, we employ the Python library ast to analyze the AST for each source code file in the code repository. From the AST, we extract six essential types of basic information: 1) Comment. 2) Signature. This knowledge illustrates the parameters and returns of a function. For example, the signature of the function _construct_text is def _construct_text(tag_elem, include_tail_text)-$>$str. 3) Function Source code. 4) Class associated with the function (if a function is a member of a class). 5) Is_Empty tag. True means empty function, False means empty label. For example, function _construct_text has is_Empty value False. 6) The file path of the function. This reflects the relative location of the current function in the entire code repository.
Semantic Code Knowledge Extraction. AST is criticized for its prone to long-term dependency problems. Also, simply employing AST potential weakens the capability to capture real and complex semantics [42]. Thus, we extract three types of semantic code knowledge to enhance the basic code knowledge, as follows:
1) Fully Qualified Name (FQN) of Functions. FQN indicates the function’s origin in the code repository and, if applicable, the specific class within the module. For example, in the local module of the Fig. 1, the function _construct_text (highlighted in yellow) has the FQN unstructured.documents.html._construct_text.
@@ III-A2 Third-Party Library Extraction
Third-party libraries often appear in code files, and when specific libraries are provided to a model, it likely has learned how to use them during training [43, 44]. Therefore, supplying just the library name allows the model to automatically understand how to use it for a given task. For example, we can ask the model to use the NumPy library to compute the cosine similarity, given the input vector_a and vector_b. In this case, the model can directly call functions in the NumPy library, such as numpy.dot and numpy.linalg.norm, to perform the required computations. The pre-trained model’s prior knowledge and extensive large-scale training enable them to effectively utilize third-party libraries, improving the efficiency and accuracy of code generation. Therefore, we extract the import statements of third-party libraries directly from all code files via AST and saved their library names to construct the third-party library base.
@@ III-B Three Types of Repo-Aware Knowledge Retrieval
As we discussed in Section II-A, although the maximum input length grows greatly, it is still impossible to feed the entire code repo and third-party library information to LLMs. Furthermore, irrelevant context not only affects model efficiency and increases expenses, but also potentially compromises the model performance [34].
To enable the model to fully utilize the three types of information extracted from the code repository, we conduct comprehensive knowledge mining: extracting four types of local information (Section III-B1), retrieving global functions from the function base using developer requirements to obtain global-aware information (Section III-B2), and structuring third-party library information from the third-party library base (Section III-B3).
@@ III-B2 Global-Module Knowledge Mining
  • Global Function Retrieval Unit: This unit uses function descriptions and the what-if function-code as input to retrieve reusable global functions from the function base. Specifically, we first embed function_description into $vector_{des}$ and compute the cosine similarity between this vector and the summary vectors of all functions in the function base. Thus, we can rank and obtain the top k $RetrievedFun_{des}$. Second, we embed the what-if function-code as a $vector_{f}$ and compute the cosine similarity between this vector and the code vectors of all the functions in the function base, so we can rank and obtain the top k $RetrievedFun_{code}$. Finally, we merge the sets of $RetrievedFun_{des}$ and $RetrievedFun_{des}$ to obtain the global function that is most relevant to the provided input, which is denoted as $RetrievedFun$.
@@ III-B3 Third-Party Library Knowledge Mining
In the process of code generation, to behavior similar to human developers, it’s crucial for the LLM to be aware of the third-party libraries pre-installed in the current code repository environment. Without this awareness, there’s a risk of generating code that relies on third-party libraries unsupported by the current environment, rendering the code non-functional. To ensure that, we create a comprehensive list of all the third-party libraries found in the third-party library base, as extracted in section III-A2. It’s important to note that we extract only the names of these libraries and do not include any of their code. Third-party libraries are publicly accessible and widely applied, so according to our empirical observation, we found that the model has already encountered and learned most of the libraries during its training process. Thus, we only extract the names of these libraries without any of their code. By constructing this list of third-party libraries, we empower the LLM to generate code that aligns with the available environment and utilizes the appropriate libraries effectively, reducing the risk of non-executable code.
@@ III-C $A^{3}$ode Generation
After section III-B, we obtain structured local-aware knowledge, global-aware knowledge, and third-party-library-aware knowledge. To ensure the model’s full awareness of these three types of knowledge, we design an $\mathbf{A^{3}}$ Code Generation Unit that seamlessly integrates user requirements with the three knowledge sources to generate $A^{3}$ functional code.
The $A^{3}$ Code Generation Unit, as depicted in Fig. 3, consists of four key components: task description, examples, input, and output. In the task description (purple part), we employ a structured prompt to delineate the repository-aware code generation task. Key components of the prompt include: 1) @persona: Guiding the model to assume the role of an expert developer, providing clarity on the task. 2) @terminology: Defining crucial information in a keyword format. 3) @instruction: This section furnishes explicit instructions for executing the task:
@@
In the @command, we guide the model through the code generation process step by step by employing chain-of-thought [45, 46, 47].
@@
In the @rule, we establish three rules to govern the code generated by the model, including calling import FQN, restricting the use of specified third-party libraries, and ensuring that the generated code is not empty (as the model may occasionally generate abstract functions). We then provide three representative examples. Finally, we input the developer requirement and our 3 knowledge into Unit to get the $A^{3}$ function code (i.e., _to_be_generated_code in Fig. 3 (b)). Notably, developers only needs to input requirements, and all 3 types of repo-aware knowledge are obtained automatically.
Due to the page limitation, we have organized all prompts for our AI units (e.g., Code Summarizer, and $A^{3}$ Code Generator) in the “prompts/” folder of our artifact [19].

### 05 Mar 2024 v4

@@ I Introduction
In the realm of Software Engineering, the advent of Large Language Models (LLMs) has revolutionized various tasks, including code generationĀ [1, 2, 3, 4, 5]. The automatic generation of code by these models significantly reduces development time and effortĀ [6]. Specifically, in the process of evolving software projects, software engineering developers are frequently tasked to add new features and functionalityĀ [7]. LLMs such as ChatGPTĀ [8], GitHub CopilotĀ [9], CodexĀ [10] are capable of automatically generating code based on the developers’ requirements (e.g., natural language descriptions of functions and function definitions).
A significant gap still persists between the code generated by LLMs and that crafted by human developers. When adding new functionality to a specific module, developers already had sufficient knowledge about the current task and the entire code repository. Also, according to a previous studyĀ [11], developers often expect that code generation tools could be aware of more context/knowledge by using specific frameworks/librariesĀ [12]. In addition, they hope the generated code maximum utilizes specific third-party libraries (e.g., invocation of an API in a library)Ā [13, 14, 15]. Above all, in contrast to straightforward LLM-based code generation, developers especially consider three primary categories of information:
@@
Global Information: This category involves functions from other code files within the same repository, accessible via fully qualified import API names. A previous research highlighted that only around 30% of methods operate independently in open-source projectsĀ [16]. Hence, these external functions are essential for implementing new functions without duplicating code. For example, developers import functions ā€œfrom unstructured.partition.text_type import is_bulleted_textā€ from other modules, instead of writing a function to judge ā€œbulleted textā€ from scratch.
@@
Unfortunately, none of existing code generation methods take above three information into consideration simultaneously. As the most common code generation tool, ChatGPTĀ [8] demonstrates commendable coding abilityĀ [17] but relies solely on user input. GitHub CopilotĀ [9], which explicitly utilizes local information spanning the current working context, often struggles to harness the full potential of global information presented in the entire code repository [^1] https://githubnext.com/projects/copilot-view/. When these methods disregard local and global module information, they risk omitting crucial logic during code generationĀ [18]. For instance, they might forget to extract text from a tag before attempting to identify it as a bullet point (is_bulleted_text). Furthermore, neglecting third-party library information may cause the LLMs to randomly select third-party libraries and generate code based on them. If the selected libraries third-party are unsupported under the given environment, it can lead to compatibility issues, such as ā€œModuleNotFoundError: No module named xxx.ā€
[…]
@@
[…]
@@
    We open-source our replication packageĀ [19], including the dataset, source code, and experimental results, for follow-up research.
@@ II-A Limitation of Existing Code Generation Approaches
Benefiting from the rapid development of Transformer structures and pre-trained LLMs in the early 2020s, code-LLMs experienced significant growth. Transformer code models, exemplified by milestone works such as CodeT5Ā [20], CodeBERTĀ [21], and GraphCodeBERTĀ [22], dominated the scene. More recently, decoder-only Transformer models have become the mainstream in generative models. Represented by CodexĀ [23], CodeGenĀ [24], and others likeĀ [25, 26, 27, 28, 29].
Nowadays, to Software Engineering developers, LLM-empowered code generation has become an irreplaceable assistant tool in their development cycleĀ [30, 31]. Although the code-LLM has been developed for years, those code generation tools still play a supporting role in the actual SE development. Essentially, there are significant differences between the code generated by LLMs and that crafted by human developers. Specifically, we conclude the phenomenon in the following threefold:
@@
    Discussion 1: Providing longer prompts with richer content seems a plausible solution. With the growth of LLM size, the maximum input length increases from 512 tokens for the early modelsĀ [32] to 32,000 tokens for the extended GPT-4 modelsĀ [33]. However, it is still impossible to feed the entire code repository and all other relevant information into the LLMs. Furthermore, studies have demonstrated that irrelevant context in prompts not only affects model efficiency and increases expenses, but also potentially compromises the model performanceĀ [34]. Therefore, to maintain the practicability and effectiveness simultaneously for code generation tasks, external context information should be extracted and selectively fed into LLMs.
    Observation 2: Existing methods focus on generating independent and often small-scale code unitsĀ [35, 16]. According to a previous studyĀ [35], the code generated by LLMs shows lower method-level correctness. Although various benchmarks have been proposed (Details are discussed in the Related Work (SectionĀ VII)), newly generated code still remains questionable.
@@
    Observation 3: Some code-LLMs (e.g., GPT-4) have developed superior skills in simple algorithms such as greedy algorithm and searching strategies, but they perform poorly in complicated data structures such as tree structure and graph theoryĀ [36].
@@ II-B Code Generation in Code Repository
During software development, programmers typically work within a specific code environment, essentially a designated code repository. This repository encompasses all project modules, functions in these modules, and third-party libraries, serving as the foundational framework for building and expanding software applications. In this context, developers, especially experienced ones, must thoroughly navigate and understand this code repository environment to ensure seamless integration of newly added code with the existing one. When it comes to LLM-based code generation techniques, solely relying on inputting developer requirements to these models fails to fully utilize the potential of LLMs in interacting with existing code repository environments. Specifically, these techniques are blind to the repository’s unique naming conventions and programming styles, particularly when the completion target is relevant to cross-file contextĀ [37, 38, 39]. Instead of using a single code file, well-designed programs emphasize separating complicated functionality into independent modules, each implementing only one aspect of the desired functionalityĀ [40, 41].
@@
Furthermore, since the model is unaware of the existence of the ā€œcheck text for bullet charactersā€ function in the global module, it defines a new function is_bullet_text (highlighted in blue) in the generated code. A large part of this code duplicates the function is_bullet_text in the global module unstructured/partition/text_type.py, resulting in code redundancy. It also brings additional concerns about the correctness of the newly generated function.
Additionally, since the model is unaware of which third-party libraries are available, it arbitrarily imports and uses the bs4 (BeautifulSoup4) library. This can result in the generated code that refers to a library that is not available, leading to the bug ā€œModuleNotFoundError: No module named ***ā€.
@@ II-C Bridging the Gap: Making LLMs Aware of Three Types of Repository Knowledge
As shown in Fig.Ā 1 Local Module Aware & Global Module Aware & Third-Party Library Aware Code Generation, when the developer requirement is entered, since the model is aware of the knowledge of the local module, the generated code reuses the _construct_text function (highlighted in yellow) in the local module to extract the text content in the text label element. In addition, because the model is aware of knowledge of the global module, the generated code reuses the is_bulleted_text function (highlighted in blue) from the global module unstructured/partition/text_type.py, thus reducing redundant code definitions. In addition, the model is aware of the existence of third-party libraries and therefore reuses the lxml library to ensure its availability.
@@ III $A^{3}$odGen Framework
In this section, we elaborate on our novel $A^{3}$-CodGen Framework to harness the full potential of code repository knowledge for assisting LLMs in code generation. As depicted in Fig.Ā 2, the $A^{3}$-CodGen Framework comprises three essential phases. First, we obtain the function base and third-party library base (SectionĀ III-A), providing sufficient knowledge to support subsequent retrieval. Then, we retrieve the code repository knowledge from other code files, the third-party library repository, and the current file. After that, we organize them into 3 types of code repository aware knowledge (SectionĀ III-B). Finally, the three code repository-aware knowledge are fused into a prompt (SectionĀ III-C), serving as input to the foundation LLM, to ultimately generate codes with three essential awareness.
@@ III-A Repository Knowledge Base Construction
Code repositories contain a wealth of knowledge to exploit, but this knowledge usually exists in an unstructured form and cannot be directly utilized effectively, scattering in the code repository. Therefore, we need to perform systematic knowledge extraction to transform and compress this knowledge into a structured form in order to build a knowledge base to support subsequent knowledge utilization, such as knowledge retrieval. Thus, we perform code repository information extraction (SectionĀ III-A1) and third-party library extraction (SectionĀ III-A2) to obtain the function base and third-party-library base, respectively.
@@ III-A1 Code Repository Information Extraction
Basic Code Knowledge Extraction. Abstract Syntax Tree (AST) is a tree representation of the source code, and studies have demonstrated its usefulness on various SE tasksĀ [42]. In this study, we employ the Python library ast to analyze the AST for each source code file in the code repository. From the AST, we extract six essential types of basic information: 1) Comment. 2) Signature. This knowledge illustrates the parameters and returns of a function. For example, the signature of the function _construct_text is def _construct_text(tag_elem, include_tail_text)-$>$str. 3) Function Source code. 4) Class associated with the function (if a function is a member of a class). 5) Is_Empty tag. True means empty function, False means empty label. For example, function _construct_text has is_Empty value False. 6) The file path of the function. This reflects the relative location of the current function in the entire code repository.
Semantic Code Knowledge Extraction. AST is criticized for its prone to long-term dependency problems. Also, simply employing AST potential weakens the capability to capture real and complex semanticsĀ [42]. Thus, we extract three types of semantic code knowledge to enhance the basic code knowledge, as follows:
1) Fully Qualified Name (FQN) of Functions. FQN indicates the function’s origin in the code repository and, if applicable, the specific class within the module. For example, in the local module of the Fig.Ā 1, the function _construct_text (highlighted in yellow) has the FQN unstructured.documents.html._construct_text.
@@ III-A2 Third-Party Library Extraction
Third-party libraries often appear in code files, and when specific libraries are provided to a model, it likely has learned how to use them during trainingĀ [43, 44]. Therefore, supplying just the library name allows the model to automatically understand how to use it for a given task. For example, we can ask the model to use the NumPy library to compute the cosine similarity, given the input vector_a and vector_b. In this case, the model can directly call functions in the NumPy library, such as numpy.dot and numpy.linalg.norm, to perform the required computations. The pre-trained model’s prior knowledge and extensive large-scale training enable them to effectively utilize third-party libraries, improving the efficiency and accuracy of code generation. Therefore, we extract the import statements of third-party libraries directly from all code files via AST and saved their library names to construct the third-party library base.
@@ III-B Three Types of Repo-Aware Knowledge Retrieval
As we discussed in SectionĀ II-A, although the maximum input length grows greatly, it is still impossible to feed the entire code repo and third-party library information to LLMs. Furthermore, irrelevant context not only affects model efficiency and increases expenses, but also potentially compromises the model performanceĀ [34].
To enable the model to fully utilize the three types of information extracted from the code repository, we conduct comprehensive knowledge mining: extracting four types of local information (SectionĀ III-B1), retrieving global functions from the function base using developer requirements to obtain global-aware information (SectionĀ III-B2), and structuring third-party library information from the third-party library base (SectionĀ III-B3).
@@ III-B2 Global-Module Knowledge Mining
[…]
@@ III-B3 Third-Party Library Knowledge Mining
In the process of code generation, to behavior similar to human developers, it’s crucial for the LLM to be aware of the third-party libraries pre-installed in the current code repository environment. Without this awareness, there’s a risk of generating code that relies on third-party libraries unsupported by the current environment, rendering the code non-functional. To ensure that, we create a comprehensive list of all the third-party libraries found in the third-party library base, as extracted in sectionĀ III-A2. It’s important to note that we extract only the names of these libraries and do not include any of their code. Third-party libraries are publicly accessible and widely applied, so according to our empirical observation, we found that the model has already encountered and learned most of the libraries during its training process. Thus, we only extract the names of these libraries without any of their code. By constructing this list of third-party libraries, we empower the LLM to generate code that aligns with the available environment and utilizes the appropriate libraries effectively, reducing the risk of non-executable code.
@@ III-C $A^{3}$ode Generation
[…]
The $A^{3}$ Code Generation Unit, as depicted in Fig.Ā 3, consists of four key components: task description, examples, input, and output. In the task description (purple part), we employ a structured prompt to delineate the repository-aware code generation task. Key components of the prompt include: 1) @persona: Guiding the model to assume the role of an expert developer, providing clarity on the task. 2) @terminology: Defining crucial information in a keyword format. 3) @instruction: This section furnishes explicit instructions for executing the task:
@@
In the @command, we guide the model through the code generation process step by step by employing chain-of-thoughtĀ [45, 46, 47].
@@
In the @rule, we establish three rules to govern the code generated by the model, including calling import FQN, restricting the use of specified third-party libraries, and ensuring that the generated code is not empty (as the model may occasionally generate abstract functions). We then provide three representative examples. Finally, we input the developer requirement and our 3 knowledge into Unit to get the $A^{3}$ function code (i.e., _to_be_generated_code in Fig.Ā 3 (b)). Notably, developers only needs to input requirements, and all 3 types of repo-aware knowledge are obtained automatically.
Due to the page limitation, we have organized all prompts for our AI units (e.g., Code Summarizer, and $A^{3}$ Code Generator) in the ā€œprompts/ā€ folder of our artifactĀ [19].